### v1

@@ Appendix B Proofs for Section 3
$W^{*}=\mathbbm{1}\left(U\leq\frac{u(X)}{\sup(\text{supp}(u(X)\mid W_{0}=1))}\right)\cdot W_{0}$
where $U\perp\!\!\!\perp(X,Y(1),Y(0),W_{0})$. Then $W^{*}$ is a regular subpopulation of $W_{0}$ and $\underline{w}^{*}(X)=\mathbb{P}(W^{*}=1\mid X,W_{0}=1)=\frac{u(X)}{\sup(\text{supp}(u(X)\mid W_{0}=1))}$ since $\frac{u(X)}{\sup(\text{supp}(u(X)\mid W_{0}=1))}\in[0,1]$ with probability 1 given $W_{0}=1$. Therefore, for all $\tau_{0}(x)=c+d^{\prime}x\in\mathcal{T}_{\text{lin}}$ we have that
$\mu(\underline{w}^{*},\tau_{0})$ $=\frac{\mathbb{E}[\underline{w}^{*}(X)\tau_{0}(X)\mid W_{0}=1]}{\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1]}$
$=\frac{\mathbb{E}\left[\frac{u(X)}{\sup(\text{supp}(u(X)\mid W_{0}=1))}\tau_{0}(X)\mid W_{0}=1\right]}{\mathbb{E}\left[\frac{u(X)}{\sup(\text{supp}(u(X)\mid W_{0}=1))}\mid W_{0}=1\right]}$

### v3

@@ Appendix G Proofs for Appendix C
Now let $x_{j}\notin\mathcal{X}^{+}$. Then
$\mathbb{P}(x_{j}\notin\widehat{\mathcal{X}}^{+})$ =($=\mathbb{P}\left(\left\{\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}(X_{i}=x_{j})=0\right\}\cup\left\{\widehat{w}_{0}(x_{j})\leq c_{n}\right\}\right)$
$\geq\mathbb{P}(\widehat{w}_{0}(x_{j})\leq c_{n})=\mathbb{P}(\sqrt{n}\widehat{w}_{0}(x_{j})\leq\sqrt{n}c_{n}).$
@@
$\mathbb{P}(\widehat{\mathcal{X}}^{+}=\mathcal{X}^{+})$ $=\mathbb{P}\left(\bigcap_{x_{j}\in\mathcal{X}^{+}}\{x_{j}\in\widehat{\mathcal{X}}^{+}\}\cap\bigcap_{x_{j}\notin\mathcal{X}^{+}}\{x_{j}\notin\widehat{\mathcal{X}}^{+}\}\right)$
$=1-\mathbb{P}\left(\bigcup_{x_{j}\in\mathcal{X}^{+}}\{x_{j}\notin\widehat{\mathcal{X}}^{+}\}\cup\bigcup_{x_{j}\notin\mathcal{X}^{+}}\{x_{j}\in\widehat{\mathcal{X}}^{+}\}\right)$
$\geq 1-\left(\sum_{j:x_{j}\in\mathcal{X}^{+}}\mathbb{P}(x_{j}\notin\widehat{\mathcal{X}}^{+})+\sum_{j:x_{j}\notin\mathcal{X}^{+}}\mathbb{P}(x_{j}\in\widehat{\mathcal{X}}^{+})\right)$
$\rightarrow 1-0=1.$
Thus, we obtain $\mathbb{P}\left(\max_{x\in\widehat{\mathcal{X}}^{+}}\widehat{a}(x)=\max_{x\in\mathcal{X}^{+}}\widehat{a}(x)\right)\geq\mathbb{P}(\widehat{\mathcal{X}}^{+}\in\mathcal{X}^{+})\rightarrow 1$, which yields
$\max_{i:\widehat{w}_{0}(X_{i})>c_{n}}\widehat{a}(X_{i})$ $=\max_{x\in\widehat{\mathcal{X}}^{+}}\widehat{a}(x)=\max_{x\in\mathcal{X}^{+}}\widehat{a}(x)+o_{p}(1).$
@@
Part 2: Asymptotic Distribution
We first establish the joint limiting distribution of terms (i) $\sqrt{n}(\frac{1}{n}\sum_{i=1}^{n}\widehat{a}(X_{i})\widehat{w}_{0}(X_{i})-\mathbb{E}[a(X)w_{0}(X)])$, (ii) $\sqrt{n}(\frac{1}{n}\sum_{i=1}^{n}\widehat{w}_{0}(X_{i})-\mathbb{E}[w_{0}(X_{i})])$, and (iii) $\sqrt{n}(\max_{i:\widehat{w}_{0}(X_{i})>c_{n}}\widehat{a}(X_{i})-\max_{x\in\mathcal{X}^{+}}a(x))$. The terms (i) and (ii) can be expanded as follows:
$\sqrt{n}$ $\left(\frac{1}{n}\sum_{i=1}^{n}\widehat{a}(X_{i})\widehat{w}_{0}(X_{i})-\mathbb{E}[a(X)w_{0}(X)]\right)$
$=\sqrt{n}\left(\sum_{j=1}^{K}\widehat{a}(x_{j})\widehat{w}_{0}(x_{j})\widehat{p}_{j}-\sum_{j=1}^{K}a(x_{j})w_{0}(x_{j})p_{j}\right)$
$=\sum_{j=1}^{K}w_{0}(x_{j})p_{j}\sqrt{n}(\widehat{a}(x_{j})-a(x_{j}))+\sum_{j=1}^{K}a(x_{j})p_{j}\sqrt{n}(\widehat{w}_{0}(x_{j})-w_{0}(x_{j}))$
$+\sum_{j=1}^{K}a(x_{j})w_{0}(x_{j})\sqrt{n}(\widehat{p}_{j}-p_{j}))+o_{p}(1)$ (A.1)
and
$\sqrt{n}$ $\left(\frac{1}{n}\sum_{i=1}^{n}\widehat{w}_{0}(X_{i})-\mathbb{E}[w_{0}(X)]\right)=\sqrt{n}\left(\sum_{j=1}^{K}\widehat{w}_{0}(x_{j})\widehat{p}_{j}-\sum_{j=1}^{K}w_{0}(x_{j})p_{j}\right)$
$=\sum_{j=1}^{K}\left(p_{j}\sqrt{n}(\widehat{w}_{0}(x_{j})-w_{0}(x_{j}))+w_{0}(x_{j})\sqrt{n}(\widehat{p}_{j}-p_{j})\right)+o_{p}(1).$ (A.2)
For term (iii), we use the expansion
$\sqrt{n}\left(\max_{i:\widehat{w}_{0}(X_{i})>c_{n}}\widehat{a}(X_{i})-\max_{x\in\mathcal{X}^{+}}a(x)\right)$ $=\sqrt{n}\left(\max_{i:\widehat{w}_{0}(X_{i})>c_{n}}\widehat{a}(X_{i})-\max_{x\in\mathcal{X}^{+}}\widehat{a}(x)\right)$ (A.3)
$+\sqrt{n}\left(\max_{x\in\mathcal{X}^{+}}\widehat{a}(x)-\max_{x\in\mathcal{X}^{+}}a(x)\right).$ (A.4)
The term in (A.3) is of order $o_{p}(1)$ because
$\mathbb{P}\left(\sqrt{n}\left(\max_{i:\widehat{w}_{0}(X_{i})>c_{n}}\widehat{a}(X_{i})-\max_{x\in\mathcal{X}^{+}}\widehat{a}(x)\right)=0\right)$ $=\mathbb{P}\left(\max_{x\in\widehat{\mathcal{X}}^{+}}\widehat{a}(x)=\max_{x\in\mathcal{X}^{+}}\widehat{a}(x)\right)$
$\geq\mathbb{P}(\widehat{\mathcal{X}}^{+}\in\mathcal{X}^{+})\rightarrow 1,$
as shown earlier.
The term in (A.4) can be analyzed using Theorem 2.1 in Fang and Santos (2019), which generalizes the delta method to the class of Hadamard directionally differentiable functions. Using Lemma G.1, we have that
$\sqrt{n}\left(\max_{x\in\mathcal{X}^{+}}\widehat{a}(x)-\max_{x\in\mathcal{X}^{+}}a(x)\right)$ $=\max_{x_{j}\in\operatorname*{arg\,max}_{x\in\mathcal{X}^{+}}a(x)}\sqrt{n}\left(\widehat{a}(x_{j})-a(x_{j})\right)+o_{p}(1)$
$=:\max_{j\in\Psi_{\mathcal{X}^{+}}}\sqrt{n}\left(\widehat{a}(x_{j})-a(x_{j})\right)+o_{p}(1).$ (A.5)
Combining the expressions in (A.1), (A.2), and (A.5) with the delta method yields